The elastic limit of steel is 8 × 108 N m-2 and its Young modulus 2 × 1011 N m-2. Find the maximum elongation of a half-meter steel wire that can be given without exceeding the elastic limit.complete answer 

Grade:12th pass

1 Answers

Khimraj
5 years ago
L ⇒ ∆L = (F/A)(L/y)\Delta y = FL/A
\DeltaL = (8*108)(0.5/2*1011)
\DeltaL = 2*10-3 m
Hope it clears............................................................................................................